### v1

# MEGC2025: Micro-Expression Grand Challenge on Spot Then Recognize and Visual Question Answering

Abstract.
Facial micro-expressions (MEs) are involuntary movements of the face that occur spontaneously when a person experiences an emotion but attempts to suppress or repress the facial expression, typically found in a high-stakes environment. In recent years, substantial advancements have been made in the areas of ME recognition, spotting, and generation. However, conventional approaches that treat spotting and recognition as separate tasks are suboptimal, particularly for analyzing long-duration videos in realistic settings. Concurrently, the emergence of multimodal large language models (MLLMs) and large vision-language models (LVLMs) offers promising new avenues for enhancing ME analysis through their powerful multimodal reasoning capabilities. The ME grand challenge (MEGC) 2025 introduces two tasks that reflect these evolving research directions: (1) ME spot-then-recognize (ME-STR), which integrates ME spotting and subsequent recognition in a unified sequential pipeline; and (2) ME visual question answering (ME-VQA), which explores ME understanding through visual question answering, leveraging MLLMs or LVLMs to address diverse question types related to MEs. All participating algorithms are required to run on this test set and submit their results on a leaderboard. More details are available at https://megc2025.github.io.
Micro-expression, Spotting, Recognition, Visual question answering
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27-October 31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer visionccs: Applied computing Psychology
Spotting Recognition STRS
TP FP FN Precision Recall F1 TP FP FN Precision Recall F1 UF1 UAR
SAMM 6 217 202 0.0269 0.0288 0.0278 1 1 1 0.1667 0.3333 0.2222 0.2222 0.1667 0.0062
CAS(ME)3 3 115 34 0.0254 0.0811 0.0387 1 1 1 0.3333 1.667 0.2222 0.2222 0.3333 0.0086
Table 1. The test set results of the ME-STR task using MEAN trained on CAS(ME)2. STRS refers to the spot-then-recognize score.
1. Introduction
When a person attempts to suppress a facial expression, typically in a high-stakes scenario, there is a possibility of an involuntary movement occurring on the face, namely a facial micro-expression (ME) (Ben et al., 2021). As such, the duration of an ME is very short, generally being no more than 500 milliseconds (ms), and is the telltale sign that distinguishes them from a normal facial expression (Li et al., 2022b). Computational analysis and automation of tasks on MEs is an emerging area in multimedia research. However, only recently, the availability of a few spontaneously induced facial ME datasets has provided the impetus to advance further from the computational aspect.
Over the years, we have organized a series of ME grand challenges (MEGCs) to explore emerging directions and advance the state of the art in ME research. To date, seven MEGCs have been held: FG’18 (Yap et al., 2018), focusing on ME recognition; FG’19 (See et al., 2019), addressing both ME spotting and recognition; FG’20 (Jingting et al., 2020), centered on ME spotting; MM’21 (Li et al., 2021), introducing ME generation and continuing with spotting; MM’22 (Li et al., 2022c), continuing with spotting and generation; MM’23 (Davison et al., 2023), with a focus on spotting; and MM’24 (See et al., 2024), which introduced the spot-then-recognize paradigm and cross-cultural spotting.
This year marks the 8th MEGC, held in conjunction with ACM Multimedia 2025. Two tasks are featured: (1) ME spot-then-recognize (ME-STR), and (2) ME visual question answering (ME-VQA).
Task 1: ME-STR. Since the surge of interest in ME research over the past decade, the community has primarily treated spotting and recognition as separate tasks. However, such separation is often impractical in real-world settings. ME recognition alone assumes that the ME sequence has already been spotted, which is rarely the case in the real world. Conversely, ME spotting without recognition fails to provide a meaningful interpretation of the emotional content. The spot-then-recognize (STR) paradigm offers a more realistic setting by first identifying the occurrence of a ME (spotting), and then classifying it into an ME category (recognition). In this task, only correctly spotted samples are passed on for emotion classification. This task is a continuation from MEGC 2024 (See et al., 2024).
Task 2: ME-VQA. The rise of multimodal large language models (MLLMs) and large vision-language models (LVLMs) has opened new opportunities for ME analysis through their advanced multimodal reasoning capabilities. In this task, traditional ME annotations, such as emotion categories and action units, are reformulated into question-answer (QA) pairs. Given an image or video sequence along with a natural language prompt, the model is expected to generate answers that describe the observed MEs, their attributes, and related information. This task introduces a novel multimodal direction for ME understanding, promoting interpretability, flexibility, and human-aligned interaction through natural language. ME-VQA is a new task introduced in MEGC 2025.
Method SAMM CAS(ME)3 Overall
Coarse Fine-Grained Coarse Fine-Grained Coarse Fine-Grained
UF1 UAR UF1 UAR BLEU ROUGE UF1 UAR UF1 UAR BLEU ROUGE UF1 UAR UF1 UAR BLEU ROUGE
Video (ZS) 0.242 0.333 0 0 0 0.295 0.262 0.333 0.057 0.048 0.089 0.315 0.256 0.333 0.057 0.048 0.065 0.312
Video (FT) 0.206 0.25 0.057 0.1 0.234 0.331 0.376 0.515 0.044 0.111 0.109 0.375 0.328 0.422 0.059 0.117 0.122 0.362
OAO (ZS) 0.267 0.333 0.000 0.000 0.031 0.332 0.438 0.444 0.000 0.000 0.038 0.322 0.366 0.400 0.000 0.000 0.036 0.328
OAO (FT) 0.300 0.417 0.111 0.167 0.182 0.366 0.401 0.455 0.054 0.143 0.021 0.323 0.396 0.461 0.061 0.143 0.027 0.339
OF (ZS) 0.3 0.333 0.19 0.3 0 0.359 0.3 0.455 0.107 0.143 0.009 0.263 0.3 0.4 0.119 0.16 0.008 0.289
OF (FT) 0.489 0.5 0.067 0.1 0 0.449 0.286 0.323 0.128 0.167 0.092 0.355 0.343 0.378 0.103 0.133 0.082 0.382
Table 2. The test set results of the ME-VQA task using Qwen2.5VL-3B trained on the curated ME-VQA dataset. The coarse emotion class refers to positive, negative, and surprise. The fine-grained emotion class refers to happiness, surprise, fear, disgust, anger, and sadness. OAO refers to the input with onset, apex, and offset frames. OF referees to the optical flow input computed between the onset and apex frames. ZS refers to zero-shot performance, while FT refers to fine-tuned performance.
2. Task 1: Spot-Then-Recognize
2.1. Task Formulation
Much work on ME analysis has focused on two separate tasks: spotting and recognizing. Only recognizing an ME class can be unrealistic in a real-world setting, as it assumes that the ME sequence has already been identified. On the other hand, the spotting task is unrealistic in its applicability since it cannot interpret the actual emotional state of the person observed. A more realistic setting, also known as spot-then-recognize (STR), performs spotting followed by recognition in a sequential manner. Only samples that have been correctly spotted in the spotting step will be passed on to the recognition step to be classified into their emotion class.
2.2. Dataset
We do not have a restriction on the training set, while we recommend participants to use SAMM-LV (Yap et al., 2020), CAS(ME)3 (Li et al., 2022b), 4DME (Li et al., 2022a), CAS(ME)2 (Qu et al., 2017), or SMIC-E-long (Tran et al., 2021) datasets for training.
For the test set of ME-STR, it contains 30 long videos, including 10 long videos from the SAMM Challenge dataset (Yap et al., 2022; Davison et al., 2018) and 20 clips cropped from different videos in CAS(ME)3 (unreleased before). The frame rate for SAMM Challenge dataset is 200fps, and the frame rate for CAS(ME)3 is 30 fps. The participants should test on this unseen dataset.
2.3. Method
We consider the micro-expression analysis network (MEAN) (Liong et al., 2023) as our baseline method. MEAN is a unified neural network architecture designed to perform both ME spotting and recognition in an end-to-end manner. It is structured with shared layers followed by two distinct branches: one for spotting and one for recognition. The spotting branch functions as a regression branch, outputting a confidence score for each frame to indicate the likelihood of an ME interval. Once the spotting branch identifies candidate intervals of MEs, the recognition branch predicts the emotion class for each interval.
2.4. Evaluation Metrics
[…]
(1) $\mathrm{STRS}=\mathrm{F1\text{-}score}_{s}\times\mathrm{F1\text{-}score}_{a}.$
The details of these metrics can be found in (Liong et al., 2023).
2.5. Experiments
Table 1 presents the quantitative results of the STR task using the MEAN network training on CAS(ME)2 and testing on the MEGC 2025 test set of ME-STR. The spotting results indicate the substantial challenge of localizing MEs in long video sequences, as evidenced by the high numbers of FPs and FNs, and the low numbers of precision, recall, and F1 in both datasets. The recognition results, evaluated on the intervals detected by the spotting module, demonstrate a slightly higher performance on certain metrics, but are ultimately constrained by the quality of the detected intervals. The overall STRS score, defined as the product of the spotting and recognition F1-scores, remains low for both datasets (0.0062 for SAMM and 0.0086 for CAS(ME)3), primarily due to the limited recall and precision achieved in the spotting stage. This outcome highlights that the bottleneck in ME analysis systems lies predominantly in the spotting task; improvements in the temporal localization of ME are likely to yield substantial gains in end-to-end analysis performance.
3. Task 2: Visual Question Answering
3.1. Task Formulation
In the ME-VQA setting, the input to the model consists of (i) a sequence of video frames capturing subtle facial movements associated with MEs, and (ii) a corresponding natural language question related to the emotions or attributes of the sequence. The model is required to generate an answer in natural language, drawing on its understanding of both the visual cues and the question of MEs. These questions can cover a wide range of attributes, from binary classification such as ”Is the action unit lip corner depressor shown on the face?” to multiclass classification like ”What is the expression class?”, and to more complex inquiries like ”What are the action units present, and based on them, what is the expression class?”.
3.2. Dataset
We do not have a restriction on the training set, while we recommend participants to use SAMM (Davison et al., 2018), CASME II (Yan et al., 2014), SMIC (Li et al., 2013), CAS(ME)3 (Li et al., 2022b), or 4DME (Li et al., 2022a) datasets for training.
We also convert the SAMM, CASME II, and SMIC annotations to a ME-VQA version [^1] The ME-VQA version dataset is available at https://megc2025.github.io/challenge.html that is suitable to train a ME-VQA model.
The ME-VQA test set contains 24 ME clips, including 7 clips from the SAMM Challenge dataset (Yap et al., 2022; Davison et al., 2018) and 17 clips cropped from different videos in CAS(ME)3 (unreleased before). The frame rate for SAMM is 200 fps and the frame rate for CAS(ME)3 is 30 fps. The participants should test on this unseen dataset.
3.3. Method
For the ME-VQA task, we employ the Qwen2.5VL-3B model (Bai et al., 2025) as our baseline method. Qwen2.5VL-3B is a recent LVLM that is based on the Qwen 2.5 architecture and incorporates a vision encoder, a language model backbone, and a cross-modal fusion module. The model is pre-trained on large-scale image-text datasets, enabling it to perform a wide range of VQA tasks with strong zero-shot and few-shot generalization capabilities.
To establish a general baseline, we consider zero-shot (ZS) and fine-tuning (FT) of the Qwen2.5VL-3B model. For fine-tuning, we applied LoRA to the vision encoder, projection layers between vision and language, as well as the query and keys of the language model. We also investigated three input types: (1) Input with equally sampled video frames. (2) Input with onset, apex, and offset frames (OAO). (3) Input with optical flow between onset and apex (OF).
3.4. Evaluation Metrics
To evaluate the performance of our ME-VQA baseline, we report metrics for both emotion classification and the overall language answer quality.
For both coarse- and fine-grained emotion classification, we use unweighted F1 Score (UF1) and Unweighted Average Recall (UAR) to ensure balanced evaluation across classes. The coarse emotion classes refer to positive, negative, and surprise. The fine-grained emotion classes refer to happiness, surprise, fear, disgust, anger, and sadness. The details of these metrics can be found in (See et al., 2019).
For all VQA answers, we report bilingual evaluation understudy (BLEU) (Papineni et al., 2002) and recall-oriented understudy for gisting evaluation (ROUGE)-1 (Chin-Yew, 2004) to assess the quality of generated text. BLEU evaluates n-gram precision between predicted and reference answers as:
(2) $\mathrm{BLEU}=\exp\left(\min\left(1-\frac{r}{c},0\right)+\sum_{n=1}^{N}w_{n}\log p_{n}\right),$
where $r$ is the reference length, $c$ is the candidate length, $p_{n}$ is the $n$-gram precision, and $w_{n}$ are the weights. The ROUGE-1 score is defined as the recall of unigram overlaps between the candidate answer $C$ and the reference answer $R$:
(3) $\mathrm{ROUGE\text{-}1}=\frac{\sum_{w\in V}\min\big{(}\mathrm{N}_{p}(w),\ \mathrm{N}_{r}(w)\big{)}}{\sum_{w\in V}\mathrm{N}_{r}(w)},$
where $V$ is the set of unique words in the reference answer, and $\mathrm{N}_{C}(w)$ and $\mathrm{N}_{R}(w)$ are the occurrences of word $w$ in the candidate and reference answers, respectively.
3.5. Experiments
The experimental results presented in Table 2 provide baseline evaluations using Qwen2.5VL-3B on the test set of the MEGC 2025 ME-VQA task.
From the results, it is evident that the smallest 3B version of Qwen2.5-VL can perform some basic ME recognition. Fine-tuning improves performance across most metrics compared to zero-shot evaluation, but the improvement is not significant, likely due to only a small number of parameters were updated using LoRA.
For expressions, especially the coarse ones, OAO (onset-apex-offset) and OF (optical flow) inputs consistently outperform the video-input results, both in zero-shot and fine-tuned conditions. This indicates that providing explicit temporal structure or motion information enables the model to better capture the subtle dynamics inherent in ME recognition. However, video-input models have better overall answer quality in terms of BLEU and ROUGE-1.